OpenStudy (anonymous):

Find the distance for the following two points A (-10,-15) (-5, -3).

OpenStudy (anonymous):

\[\sqrt{49+100}\] 12.2 is Answer

OpenStudy (anonymous):

(-10)-(-5)=5 (-15)-(-3)=12 25+144=169 sqrt of 169=13

OpenStudy (anonymous):

final answer is 13

OpenStudy (anonymous):

Oops sorry \[\sqrt{144+25}=\sqrt{169}=13\]